OpenStudy (anonymous):

Is sqrt(3)(cos(225 degrees) + isin(225 degrees)) the polar form of 3-3i?

OpenStudy (anonymous):

\[\sqrt{3}(\cos(225 degrees) + i \sin(225 degrees))\]

OpenStudy (jdoe0001):

how did you find the modulus?

OpenStudy (anonymous):

Wait, I know what I did wrong. I thought it was modulus(r) = \[\sqrt{3^{2} - 3^{2}}\], but it's \[\sqrt{3^{2} + (-3)^{2}}\].

OpenStudy (jdoe0001):

modulus is \(\sqrt{3^{2} + 3^{2}}\)

OpenStudy (anonymous):

the question is not about finding anything other than \(\cos(225)\) and \(\sin(225)\) then multiply

OpenStudy (jdoe0001):

so \(\large \sqrt{18} = \sqrt{3^2 \times 2} \)

OpenStudy (jdoe0001):

your \(tan(\theta) = \frac{3}{3} = 1\)

OpenStudy (anonymous):

I thought \[\tan(\Theta) = (-3)/3\] since b = -3.

OpenStudy (jdoe0001):

\(tan^{-1}{(1)} = 45^o\) on the I quadrant, because the "x" and "y" are both positive

OpenStudy (jdoe0001):

ohh, yes... it's ... -3 shoot

OpenStudy (anonymous):

That's fine, thanks for helping.

OpenStudy (jdoe0001):

ok, then .. the modulus is the same (-3)^2 is 3 anyway, the angle will just be on the IV quaddrant, "x" is positive, "y" negative

OpenStudy (anonymous):

I got 315 degrees in QIV.

OpenStudy (jdoe0001):

yes